Show that the inverse of the function x^2 - 4x + 3 is 2 + sqrt(1+x)
Answer by lwsshak3(11628)

Show that the inverse of the function
x^2 - 4x + 3 is 2 + sqrt(1+x)
**
y=x^2-4x+3
to find the inverse, interchange x and y, then solve for y.
x=y^2-4y+3
y^2-4y=x-3
complete the square
y^2-4y+4=x-3+4
(y-2)^2=x+1
y-2=√(x+1)
y^-1=2+√(1+x)